(*) If x =a (sec theta + tan theta) square, y = b (sec theta – tan theta) square then x square.y square =

(*)
If x =a (sec theta + tan theta) square, y = b (sec theta – tan theta) square then x square.y square =

Grade:11

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below

x =a (sec theta + tan theta)2
y = b (sec theta – tan theta)2

x*y =a (sec theta + tan theta)2 * b (sec theta – tan theta)2
= ab (sec2 theta – tan 2theta)2
= ab [because sec2 theta – tan 2theta = 1]

therefore x2 * y2 = 1